# Distributional Shrinkage I: Universal Denoisers in Multi-Dimensions

### v1

@@ Abstract
Keywords— Denoising equation, Bayes shrinkage, distributional shrinkage, optimal transport, Monge-Ampère equation, generalized moment matching.
@@ 1 Introduction
We revisit the classic denoising problem in multi-dimensions with $d\in\mathbb{N}^{+}$. Let $X,Z\in\mathbb{R}^{d}$ be two independent random variables, where $X$ is the signal of interest drawn from an unknown distribution $P_{X}$, and $Z$ is the noise with distribution $P_{Z}$ that is symmetric around zero and also unknown. We only observe the noisy measurement $Y$, with an additive noise level $\sigma\in\mathbb{R}_{>0}$ known a priori:
@@
These denoisers achieve $O(\sigma^{4})$ and $O(\sigma^{6})$ accuracy in distributional matching, respectively, representing an order-of-magnitude improvement over the classical Bayes-optimal shrinkage $\mathbf{T}^{*}(y)$. These denoisers are universal, meaning they work under a host of smooth signal distributions $P_{X}$ and noise distributions $P_{Z}$ that satisfy mild moment conditions. These statements will be made precise in the following sections.
@@
This over-shrinkage effect can be explained as follows: this data-level shrinkage focuses solely on minimizing the MSE, a (conditional) first-moment matching criterion. It does not aim to match higher-order moments of the distribution. Note that the trivial map $\mathbf{T}^{\mathrm{id}}(y)=y$ matches the first moment exactly but overshoots the second moment by $\Theta(\sigma^{2})$:
@@
Figure 2: Comparison of shrinkage maps for $d=2$. We compare four denoisers as in Figure 1, but with scatterplots. Here, the signal $X$ is Gaussian, and the noise parameter $\eta:=\sigma^{2}/2$ is chosen as 0.5. We report the Wasserstein and energy distances between the denoised and true signal distributions in the top-left corner of each subplot. Again, the Bayes denoiser shrinks the distribution too aggressively. The proposed denoisers $\mathbf{T}_{1}^{f_{*}}$ and $\mathbf{T}_{2}^{f_{*},g_{*}}$ achieve improvement in distributional matching.
@@
The paper is organized as follows. In Section 2, we discuss the relations to the existing literature. In Section 3.1, we introduce the differential equations characterizing the optimal denoisers along with practical implementations, followed by assumptions on signal and noise distributions in Section 3.2. In Section 4, we study the theoretical performance of the proposed denoisers in terms of generalized moment matching. In Section 5, we establish the optimality of the proposed denoisers in terms of approximating the static Monge-Ampère equation with higher-order accuracy. In Section 6, we empirically evaluate the optimal first- and second-order denoisers and compare them with the Bayes optimal denoisers. The proofs of the main theorems are presented in Section A, while technical lemmas and their proofs are included in Section B.
@@ 2 Relations to the Literature
We work in a regime with a constant noise level, where $\sigma<1$ is a known constant, but the exact noise distribution is unknown. In this case, exact deconvolution is intractable due to the lack of knowledge of the noise distribution. We design simple, universal denoising maps that depend only on the score function of the noisy measurement distribution $q=P_{Y}$. Our denoisers achieve $O(\sigma^{4})$ and $O(\sigma^{6})$ accuracy in distributional matching for a range of signal distribution $P_{X}$ and noise distribution $P_{Z}$, respectively, improving over the classical Bayes-optimal shrinkage.
Optimal transport theory [Vil08, AGS05] offers a new framework for comparing and transforming probability distributions. Taking an optimal transport perspective also opens new avenues for revisiting the denoising problem. Due to the variational formulation of the Fokker–Planck equation [JKO98], adding small Gaussian noise to a distribution can be interpreted as gradient descent (with a small step size) in the space of probability measures with the entropy functional as the objective and Wasserstein distance as the geometry. Therefore, denoising can be viewed as the reverse process of this gradient descent; see Proposition 1 in [LDK24]. As a result, we can study denoising from the perspective of transport maps. Recent works have explored the use of optimal transport maps for denoising [Kim+23, Che+23, LDK24] and generative modeling [HGL24, LDK24, DL25].
Our approach leverages ideas from optimal transport to design denoisers that achieve distributional matching with higher accuracy. The accuracy is measured either by integral probability metrics [Mül97, Gre+12, Lia19, Lia21] induced by smooth test functions, and by approximations to the static Monge-Ampère equation [Caf92, DL25]. The set of smooth test functions can be interpreted as generalized moments of the distribution [Han82]. The static Monge-Ampère equation characterizes the optimal transport map between two distributions [Caf92, Vil08]. This work builds on these ideas to derive differential equations characterizing the optimal denoisers that achieve higher-order distributional matching.
@@ 4.1 First-Order Denoiser
Some explanations are in order. First, the integrability conditions impose decay conditions on $m$ and its gradient $\nabla m$ at infinity, which further implies that $m$ and $\nabla m$ vanish at infinity. This rules out unbounded test functions that grow to infinity, such as polynomials. However, there is a host of test functions that satisfy these conditions, for example, compactly supported smooth functions. Second, when the signal distribution $P_{X}$ has a compact support or decays fast at infinity, it is natural only to consider test functions supported on a compact set $\Omega\subset\mathbb{R}^{d}$, with smooth extensions to the entire space $\mathbb{R}^{d}$ that vanish. Such a consideration is without loss of generality for bounded test functions, or unbounded test functions with a slow growth compared to the decay of $P_{X}$ at infinity. Third, let $\varepsilon>0$, one can consider $m$ to be any continuous function supported on a compact set $\Omega$ (including polynomials), then consider smoothed test functions $m*\varphi_{\varepsilon}$ convolved with a mollifier $\varphi_{\epsilon}$ [Eva22], say $\varphi_{\varepsilon}(x)=\varepsilon^{-d}\varphi(x/\varepsilon)$ with
@@
This theorem states that for a smooth test function $m\in\mathcal{M}^{2}{\mathbb{R}^{d}}$, the first-order denoiser $\mathbf{T}_{1}^{f_{*}}$ matches the generalized moments (induced by $m$) of the signal distribution $P_{X}$ up to an error of order $O(\sigma^{4})$. Such a result holds as long as the signal distribution $P_{X}$ is smooth enough (up to 4-th order), and the noise distribution $P_{Z}$ has bounded 4-th moments and is uncorrelated. In this sense, the first-order denoiser is universal to a broad class of signal and noise distributions; it is agnostic to exact forms of both $P_{X}$ and $P_{Z}$, allowing both to be general nonparametric distributions. Thus, we call them universal denoisers.
@@ A.1 Proofs for Section 4
if (i) $\int_{\mathbb{R}^{d}}|m(y)|\differential y<\infty, d$$\int_{\mathbb{R}^{d}}|\partial_{i}m(y)|\differential y<\infty$, and (ii) $m\nabla p$ vanishes as $y\rightarrow\infty$ as $m$ vanishes and $\nabla p$ is bounded due to Assumption 1. To be more specific, (ii) is a direct application of integration by parts
@@
[…]
@@
[…]

### v2

@@ Abstract
Keywords— Universal denoisers, denoising equations, distributional shrinkage, optimal transport, Monge-Ampère equation, generalized moment matching.
@@ 1 Introduction
We revisit the classic denoising problem in multi-dimensions with $d\in\mathbb{N}^{+}$. Let $X,Z\in\mathbb{R}^{d}$ be two independent random variables, where $X$ is the signal of interest drawn from an unknown distribution $P_{X}$, and $Z$ is the noise with distribution $P_{Z}$ that is symmetric around zero and also unknown. We only observe the noisy measurement $Y$, with an additive noise level $\sigma\in(0,1)$ known a priori:
@@
These denoisers achieve $O(\sigma^{4})$ and $O(\sigma^{6})$ accuracy in distributional matching, respectively, representing an order-of-magnitude improvement over the classical Bayes-optimal shrinkage $\mathbf{T}^{*}(y)$. These denoisers are universal denoisers, meaning they work under a host of smooth signal distributions $P_{X}$ and noise distributions $P_{Z}$ that satisfy mild moment conditions. These statements will be made precise in the following sections.
@@
This over-shrinkage effect is known [GS24, JIS25] and can be explained as follows: this data-level shrinkage focuses solely on minimizing the MSE, a (conditional) first-moment matching criterion. It does not aim to match higher-order moments of the distribution. Note that the trivial map $\mathbf{T}^{\mathrm{id}}(y)=y$ matches the first moment exactly but overshoots the second moment by $\Theta(\sigma^{2})$:
@@
Figure 2: Comparison of shrinkage maps for $d=2$. We compare four denoisers as in Figure 1, but with scatterplots. Here, the signal $X$ is Gaussian, and the noise parameter $\eta:=\sigma^{2}/2$ is chosen as 0.5. We report the Wasserstein distance between the denoised and true signal distributions in the top-left corner of each subplot. Again, the Bayes denoiser shrinks the distribution too aggressively. The proposed denoisers $\mathbf{T}_{1}^{f_{*}}$ and $\mathbf{T}_{2}^{f_{*},g_{*}}$ achieve improvement in distributional matching.
@@
On the Noise Level
We remark that $\sigma\in(0,1)$ can be assumed without loss of generality. If $\sigma\geq 1$, one can rescale $Y$ by $\tau:=\|\mathop{\mathbb{E}}[YY^{\top}]\|_{\mathrm{op}}^{1/2}>\sigma$ to ensure that the effective noise level $\widetilde{\sigma}:=\sigma/\tau<1$, as long as $X$ is not a point mass at 0; rescaling $Y$ by $1.1\sigma$ also suffices as $\widetilde{\sigma}=1/1.1$. Therefore, throughout the paper, we assume $\sigma\in(0,1)$ and thus $\eta=\sigma^{2}/2\in(0,1/2)$ without loss of generality.
@@
The paper is organized as follows. In Section 2, we discuss the relations to the existing literature. In Section 3.1, we introduce the differential equations characterizing the optimal denoisers along with practical implementations, followed by assumptions on signal and noise distributions in Section 3.2. In Section 4, we study the theoretical performance of the proposed denoisers in terms of generalized moment matching. In Section 5, we establish the optimality of the proposed denoisers in terms of approximating the static Monge-Ampère equation with higher-order accuracy. In Section 6, we empirically evaluate the optimal first- and second-order denoisers and compare them with the Bayes-optimal denoisers. The proofs of the main theorems are presented in Section A, while technical lemmas and their proofs are included in Section B.
@@ 2 Relations to the Literature
We work in a regime with a constant noise level, where $\sigma$ is a known constant, but the exact noise distribution is unknown. In this case, exact deconvolution is intractable due to the lack of knowledge of the noise distribution. We design simple, universal denoising maps that depend only on the score function of the noisy measurement distribution $q=P_{Y}$. Our denoisers achieve $O(\sigma^{4})$ and $O(\sigma^{6})$ accuracy in distributional matching for a range of signal distribution $P_{X}$ and noise distribution $P_{Z}$, respectively, improving over the classical Bayes-optimal shrinkage.
Optimal transport theory [Vil08, AGS05] offers a new framework for comparing and transforming probability distributions. Taking an optimal transport perspective also opens new avenues for revisiting the denoising problem. Due to the variational formulation of the Fokker–Planck equation [JKO98], adding small Gaussian noise to a distribution can be interpreted as gradient descent (with a small step size) in the space of probability measures with the entropy functional as the objective and Wasserstein distance as the geometry. Therefore, denoising can be viewed as the reverse process of this gradient descent; see Proposition 1 in [LDK24]. As a result, we can study denoising from the perspective of transport maps. Recent works have explored the use of optimal transport maps for denoising [Kim+23, Che+23, LDK24] and generative modeling [HGL24, LDK24, DL25]. Closest to ours are [GS24, JIS25, Lia25]. [GS24, JIS25] also notice the over-shrinkage effect of the empirical Bayes denoiser. They suggest post-processing empirical Bayes methods to correct for distributional matching, assuming knowledge of $P_{X}$ or its moments. In contrast, our denoisers are universal and do not require any knowledge of $P_{X}$. [Lia25] studies distributional shrinkage under Gaussian noise in a one-dimensional setting but derives a hierarchy of denoisers that leverage higher-order score functions. In contrast, our work studies the problem in a multi-dimensional setting and constructs first- and second-order denoisers that remain universal to a broad range of noise distributions beyond Gaussian.
Our approach leverages ideas from optimal transport to design denoisers that achieve distributional matching with higher accuracy. The accuracy is measured either by integral probability metrics [Mül97, Gre+12, Lia19, Lia21] induced by smooth test functions, and by approximations to the static Monge-Ampère equation [Caf92, DL25]. The set of smooth test functions can be interpreted as generalized moments of the distribution [Han82]. The static Monge-Ampère equation characterizes the optimal transport map between two distributions [Caf92, Vil08]. Our work builds on these ideas to derive differential equations characterizing the optimal denoisers that achieve higher-order distributional matching.
@@ 4.1 First-Order Denoiser
Some explanations are in order. First, the integrability conditions impose decay conditions on $m$ and its gradient $\nabla m$ at infinity. This rules out unbounded test functions that grow to infinity, such as polynomials. However, there is a host of test functions that satisfy these conditions, for example, compactly supported smooth functions. Second, when the signal distribution $P_{X}$ has a compact support or decays fast at infinity, it is natural only to consider test functions supported on a compact set $\Omega\subset\mathbb{R}^{d}$, with smooth extensions to the entire space $\mathbb{R}^{d}$ that vanish. Such a consideration is without loss of generality for bounded test functions, or unbounded test functions with a slow growth compared to the decay of $P_{X}$ at infinity. Third, let $\varepsilon>0$, one can consider $m$ to be any continuous function supported on a compact set $\Omega$ (including polynomials), then consider smoothed test functions $m*\varphi_{\varepsilon}$ convolved with a mollifier $\varphi_{\epsilon}$ [Eva22], say $\varphi_{\varepsilon}(x)=\varepsilon^{-d}\varphi(x/\varepsilon)$ with
@@
This theorem states that for a smooth test function $m\in\mathcal{M}^{2}(\mathbb{R}^{d})$, the first-order denoiser $\mathbf{T}_{1}^{f_{*}}$ matches the generalized moments (induced by $m$) of the signal distribution $P_{X}$ up to an error of order $O(\sigma^{4})$. Such a result holds as long as the signal distribution $P_{X}$ is smooth enough (up to 4-th order), and the noise distribution $P_{Z}$ has bounded 4-th moments and is uncorrelated. In this sense, the first-order denoiser is universal to a broad class of signal and noise distributions; it is agnostic to exact forms of both $P_{X}$ and $P_{Z}$, allowing both to be general nonparametric distributions. Thus, we call them universal denoisers.
@@ 6 Numerics
Acknowledgements
The author acknowledges the support from the NSF Career Award (DMS-2042473) and the Wallman Society of Fellows at the University of Chicago. The author also thanks Nabarun Deb, Shuo-Chieh Huang, Nikos Ignatiadis, Ben Recht, and Bodhi Sen for comments on an earlier draft. This work is part I of a two-part series on distributional shrinkage; part II can be found in [Lia25].
@@ A.1 Proofs for Section 4
if (i) $\int_{\mathbb{R}^{d}}|m(y)|\differential y<\infty, d$$\int_{\mathbb{R}^{d}}|\partial_{i}m(y)|\differential y<\infty$, and (ii) $m\nabla p$ vanishes as $y\rightarrow\infty$ because $m$ vanishes and $\nabla p$ is bounded due to Assumption 1. To be more specific, (ii) is a direct application of integration by parts
@@
and the boundary condition as $m\nabla p,m\nabla\Delta p$ vanishes on the boundary as $y\rightarrow\infty$. Here we use the fact that $m$ vanishes and $\nabla p,\nabla\Delta p$ are bounded (Assumption 1). Now invoke Lemma 1(ii) again, we have element-wise the following expansions
@@
Note that $q\langle\nabla m,\nabla f\rangle\nabla f$ vanishes on the boundary as $y\rightarrow\infty$ due to the fact that $\partial_{i}m$ vanishes and $q\partial_{i}f\partial_{j}f$ is bounded (Assumption 3). Therefore, applying integration by parts and recalling (A.7), we have